OpenStudy (anonymous):

equation of the line: (0,-2) (3,2) explainations are helpful :D

OpenStudy (anonymous):

Let's find the slope, 2- -2 4 -------- = -------------- 3-0 3

OpenStudy (anonymous):

and then?....

OpenStudy (anonymous):

use the slope intercept equation of a line: y = mx + b, where m is the slope and b is the y intercept. Get the slope with change in y over change in x, or (2 - -2) / (3 - 0) = 4/3. Then plug that in for m and the x and y from either point into the equation above and solve for b: y = mx + b -2 = 0 + b b = -2 So the final equation is y = 4/3 x -2